OpenStudy (anonymous):

If f(x)=1/^3SQRTx Find F''(x)

OpenStudy (mathmale):

Are you certain you've typed the problem in as presented to you? that 1/^3 makes me nervous! Could you possibly enter your problem statement using the equation editor, below?

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad \frac{1}{\sqrt[3]{x}}\]Is this the function? Cube root in the bottom?

OpenStudy (anonymous):

Yup thats the one

OpenStudy (mathmale):

HHH: Note that if zepdrix is correct, "SQRT" does not belong in the problem statement. Instead, type "CUBE ROOT OF X."

zepdrix (zepdrix):

The equation editor uses sqrt for every root.. :) So maybe it was presented that way.

OpenStudy (mathmale):

Mybest advice at this point is to rewrite zepdrix' correct expression as y=x^(-1/3). You familiar with that?

OpenStudy (anonymous):

I don't know how to use this website much , anyways , anyone know how to find it?

OpenStudy (mathmale):

So, in the Equation Editor, \[y=\frac{ 1 }{ 3^{(1/3)} }=y=3^{(-1/3)}\]

OpenStudy (anonymous):

@mathmale I just need all the working out and steps at this point, cause fractions make me cry

OpenStudy (anonymous):

Oh yeah i got it from there , thanks man

OpenStudy (mathmale):

OK. This truly is the easiest way to solve this problem. AFter having re-written the original function, finding the 1st and 2nd derivatives should be a snap. You're welcome! :)

OpenStudy (mathmale):

Actually, I've made a mistake. Should have given you:

OpenStudy (mathmale):

\[y=x ^{-1/3}\]

OpenStudy (mathmale):

Nobody's perfect, least of all I!

OpenStudy (anonymous):

I really appreciate the time you're putting into my question.. True that, we all make mistakes, wish me luck man , few hours left to my finals, currently surviving on coffee